# Uncertainty Calibration of Multi-Label Bird Sound Classifiers

### v1

@@ Abstract
Passive acoustic monitoring enables large-scale biodiversity assessment, but reliable classification of bioacoustic sounds requires not only high accuracy but also well-calibrated uncertainty estimates to ground decision-making. In bioacoustics, calibration is challenged by overlapping vocalisations, long-tailed species distributions, and distribution shifts between training and deployment data. The calibration of multi-label deep learning classifiers within the domain of bioacoustics has not yet been assessed. We systematically benchmark the calibration of four state-of-the-art multi-label bird sound classifiers on the BirdSet benchmark, evaluating both global, per-dataset and per-class calibration using threshold-free calibration metrics (ECE, MCS) alongside discrimination metrics (cmAP). Model calibration varies significantly across datasets and classes. While Perch v2 and ConvNeXtBS show better global calibration, results vary between datasets. Both models indicate consistent underconfidence, while AudioProtoPNet and BirdMAE are mostly overconfident. Surprisingly, calibration seems to be better for less frequent classes. Using simple post hoc calibration methods we demonstrate a straightforward way to improve calibration. A small labelled calibration set is sufficient to significantly improve calibration with Platt scaling, while global calibration parameters suffer from dataset variability. Our findings highlight the importance of evaluating and improving uncertainty calibration in bioacoustic classifiers.
Figure 1: Calibration evaluation of AudioProtoPNet, BirdMAE, ConvNeXtBS and Perch v2 globally on all BirdSet test samples visualised in a reliability diagram (Section 2.1). Mean predicted probabilities are shown on the x-axis, while empirical frequencies are shown on the y-axis. The miscalibration score (MCS) is outlined for each model.
@@ 1 INTRODUCTION
Human pressures are leading to declines in biodiversity necessitating the development of non-invasive measurement systems and effective analysis tools to monitor these trends [Schmeller et al., 2017, Gosselin and Callois, 2018]. Bioacoustics describes the study of sounds produced by organisms, offering a non-invasive measurement method of ecological habitats and its species [Laiolo, 2010]. Passive acoustic monitoring (PAM)has emerged as a powerful technique for continuous and non-invasive biodiversity assessments of animal populations in remote areas through low-cost recorders at scale [Sugai et al., 2019]. Recent advances in deep learning (DL)have made it feasible to process large collections of PAM recordings and automatically detect species vocalisations, enabling applications such as population trend estimation [McGinn et al., 2023], impact assessments [Kershenbaum et al., 2025], and rapid alerts for protected or invasive species [Baumgartner et al., 2019]. While performance of DLmodels in bioacoustics has shown steady increases in previous years, generalisation to soundscape recordings with overlapping vocalisations, as collected in PAMscenarios, is still challenging [Stowell, 2022, Rauch et al., 2024, Schwinger et al., 2025]. In practice, decisions based on classifier outputs rarely stand alone: ecologists, regulators, and land managers need to understand when to rely on model outputs, and when to defer to human review or collect more data. Variability in model calibration across domains calls into question the validity of fixed thresholds when applied across domains [Navine et al., 2024]. Reliable decision-making therefore requires not only high discrimination (e.g., AUROC, cmAP scores) but also well-calibrated predictive probabilities that honestly reflect uncertainty [Guo et al., 2017].
@@
Calibration is especially challenging in bird sound classification as the task is inherently multi-label due to overlapping vocalisations. In addition, soundscapes vary widely across habitats and devices and species occurrence follow a long tail, with rare species being of greatest conservation concern [Stowell, 2022]. Distribution shift between training data (often directional focal recordings) and deployment data (omnidirectionally recorded soundscapes) can further degrade reliability [Rauch et al., 2024]. In these settings, overconfident models can bias occupancy predictions, inflate apparent detection rates, and mislead thresholding and decision making, leading to samples not flagged for human evaluation. Conversely, underconfidence reduces the utility of scores for prioritising expert review. Reliable uncertainty estimates are therefore critical to support transparent and accountable use of PAM in environmental decision-making.
@@
Despite growing interest, the calibration behaviour of multi-label audio classifiers remains underexplored. Many studies prioritise accuracy or ranking metrics, while probability quality is reported sparsely or via single-number summaries that may obscure class-specific effects [van Merriënboer et al., 2025, Kahl et al., 2021]. Moreover, commonly used post hoc calibration methods, which improve model calibration without retraining, are not uniformly evaluated in this domain. Based on this motivation, our contributions can be summarised as follows:
@@
Contributions C1. Benchmark calibration of four state-of-the-art multi-label bird sound classifiers on BirdSet, reporting global, per-dataset and per-class calibration with threshold-free metrics, revealing that per-domain (e.g. dataset, class etc) calibration is highly variable. C2. Propose post hoc strategies (temperature and Platt scaling) showing that Platt scaling with 10 minutes of test data can significantly improve calibration. C3. Summarise implications to give practitioners concrete suggestions for deployment specific calibration evaluation and improvement. Code is available at [^1] https://github.com/DBD-research-group/UncertainBird.
@@ 2.1 Uncertainty and Calibration
We follow the common decomposition into (i) aleatoric uncertainty inherent in the data (e.g., noise, wind, overlapping vocalisations, distant calls) and (ii) epistemic uncertainty due to a lack of model knowledge (e.g., rare species, novel soundscapes, or data-scarce regions) [Rewicki and Gawlikowski, 2022]. We focus on the latter and assume that acquiring more data is not possible.
@@
Calibration methods can be broadly divided into post hoc and ad hoc approaches. Post hoc methods operate on a trained model’s outputs, learning a mapping from uncalibrated logits to calibrated probabilities without modifying the model weights. Temperature scaling (TS) [Guo et al., 2017], Platt scaling (PS) [Platt et al., 1999], isotonic regression [Niculescu-Mizil and Caruana, 2005], and vector scaling [Kull et al., 2019] are widely used post hoc techniques. These methods are attractive due to their simplicity and low computational overhead. Bayesian deep learning methods, such as Monte Carlo dropout [Gal and Ghahramani, 2016], deep ensembles [Lakshminarayanan et al., 2017], and variational inference [Blundell et al., 2015, Rizos et al., 2024], provide principal uncertainty estimates but add computational overhead.
Ad hoc calibration, in contrast, integrates calibration objectives directly into model training. This includes approaches such as label smoothing [Müller et al., 2019], focal loss [Lin et al., 2017], and mixup [Zhang et al., 2018], which implicitly encourage calibrated predictions by modifying the loss. More recently, spectral-normalised neural Gaussian processes [Liu et al., 2020] have been proposed to enhance distance awareness in DL, potentially improving calibration. Ad hoc methods can yield better calibration but often require careful tuning and may increase training complexity.
@@ 2.2 Platt and Temperature Scaling
Modern bioacoustic models typically formulate animal sound detection as a multi-label classification problem [Rauch et al., 2024, van Merriënboer et al., 2025, Heinrich et al., 2025]. For each class $c\in\{1,\dots,C\}$, the model produces a logit $z_{c}\in\mathbb{R}$, which is mapped to a confidence score $\hat{p}_{c}\in[0,1]$ via the sigmoid function, i.e., $\hat{p}_{c}=\sigma(z_{c})$. The resulting confidence scores can be calibrated post hoc without retraining the base model [Platt et al., 1999, Niculescu-Mizil and Caruana, 2005, Guo et al., 2017]. A widely used method is TS [Guo et al., 2017], which rescales the logits by a positive temperature parameter $T\in\mathbb{R}_{>0}$ and then re-applies the sigmoid:
@@
The parameter $T$ controls the sharpness of the resulting distribution of calibrated confidence scores, see Figure 2(a). Setting $T=1$ recovers the uncalibrated model. For $T>1$, confidences are smoothed. Values below $0.5$ move up toward $0.5$, and values above $0.5$ move down toward $0.5$. For $T<1$, the distribution is sharpened, pushing values below $0.5$ toward 0 and values above $0.5$ toward $1$. If the same temperature parameter is used for all classes $c$, TSpreserves the monotonic ordering of the model’s predicted confidence scores [Guo et al., 2017, Tomani et al., 2022]. Because $\sigma(0)=0.5$, confidence scores exactly at $0.5$ remain unchanged under TS. As a consequence, purely systematic offset errors (e.g., when the reliability curve [Murphy and Winkler, 1977] lies consistently above or below the diagonal) cannot be corrected in this way [Alexandari et al., 2020]. If offset errors need to be corrected, PS [Platt et al., 1999] provides a useful generalisation of TS:
@@
In addition to the scaling parameter $T$, a bias $b\in\mathbb{R}$ is learned that shifts the entire sigmoid curve up or down, thereby changing the point at which a confidence score of $0.5$ is achieved (see Figures 2(b) and 2(c)). Positive biases ($b>0$) increase confidence scores, while negative biases ($b<0$) decrease them. Both TSand Platt scaling can also be parameterised on a per-class basis, using class-specific parameters $T_{c}$ and $b_{c}$ to compensate for class-dependent underconfidence and overconfidence [Alexandari et al., 2020, Frenkel and Goldberger, 2021]. In practice, the parameters are estimated by minimising the negative log-likelihood (NLL)of the true labels on a held-out validation set [Lin et al., 2007, Guo et al., 2017]. The resulting optimisation problems are convex and involve only a few parameters to learn [Lin et al., 2007, Guo et al., 2017], making them well suited even when only limited data are available for calibration [Niculescu-Mizil and Caruana, 2005]. After fitting, the learned parameters are applied to all future predictions without modifying the base model.
@@ 2.3 Related work
Uncertainty calibration in multi-label classification
Multi-label classification presents unique calibration challenges due to label imbalance and the presence of co-occurring classes. Recent research has highlighted that standard deep neural networks, when trained for multi-label tasks, often produce poorly calibrated probability estimates, especially in the presence of severe class imbalance and large label spaces. Cheng and Vasconcelos [Cheng and Vasconcelos, 2024] demonstrate that popular asymmetric loss functions, while effective for improving accuracy in unbalanced datasets, do not guarantee well-calibrated outputs. They attribute this to the lack of the strictly proper property in these losses and propose a new Strictly Proper Asymmetric loss, along with a label pair regulariser, to improve calibration without sacrificing accuracy. Their experiments show significant reductions in calibration error across diverse multi-label datasets.
In the context of extreme multi-label classification (XMLC), where the number of possible labels can reach into the millions, Ullah et al. [Ullah et al., 2024] systematically evaluate the calibration of a wide range of XMLC models. They find that naive application of standard metrics such as ECE can be misleading in long-tailed label distributions. To address this, they introduce calibration@k metrics (e.g., ECE@k), which focus on the top-k predicted labels, providing a more meaningful assessment of calibration in XMLC scenarios. Furthermore, they show that post hoc calibration using isotonic regression can substantially improve reliability without degrading predictive performance, making it a practical tool for practitioners. In bioacoustics, the use of pretrained models (e.g. Perch v2 [van Merriënboer et al., 2025], BirdNET [Kahl et al., 2021]) is increasingly common due to a lack of labelled in-domain data. At the same time, the number of classes for which these mode s predict over is increasing, with models such as Perch v2 having almost 15,000 classes available.
@@ 2.3.1 Uncertainty Calibration in Audio Classification
Uncertainty calibration for deep audio classifiers is an emerging research area, motivated by the need for reliable confidence estimates in practical applications such as multimedia retrieval, urban sound monitoring, and bioacoustic surveys. While deep neural networks have achieved state-of-the-art performance in audio classification tasks [Chen et al., 2022], model calibration is not well-investigated, yet.
Ye et al. [Ye et al., 2022] provide one of the first systematic investigations of calibration methods for deep audio classifiers. They benchmark the calibration methods Monte Carlo Dropout, deep ensembles, focal loss, and spectral-normalised neural Gaussian processes (SNGP) on environmental sound and music genre classification datasets. Their results indicate that uncalibrated models are frequently overconfident, and that SNGP in particular achieves strong calibration and computational efficiency. Notably, while focal loss can improve calibration, it may degrade out-of-distribution detection performance, highlighting trade-offs between calibration and other uncertainty-related objectives.
In the context of bioacoustics, calibration is further complicated by the scarcity of strongly labelled data and the prevalence of rare species [Stowell, 2022]. Rizos et al. [Rizos et al., 2024] propose a Bayesian variational method which propagates uncertainty (prediction and variance) using a Bayesian neural network and is used for post hoc label smoothing.
Table 1: Overview of the different datasets in BirdSet [Rauch et al., 2024]. J refers to the class imbalance index.
mainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmain Set ||Annotations|| ||Duration|| $\boldsymbol{J}$ $\boldsymbol{C}$ Train XCL 712.433  7,200h 0.85 9,734 Val POW 16,052 6.3h 0.66 48 Test PER 14,798 21h 0.78 132 NES 6,952 34h 0.76 89 UHH 59,583 50.9h 0.64 25 HSN 10,296 16.7h 0.54 21 NBP 5,493 0.8h 0.92 51 SSW 50,760 285h 0.77 81 SNE 20,147 33h 0.70 56

### v2

@@ Abstract
Passive acoustic monitoring enables large-scale biodiversity assessment, but reliable classification of bioacoustic sounds requires not only high accuracy but also well-calibrated uncertainty estimates to ground decision-making. In bioacoustics, calibration is challenged by overlapping vocalisations, long-tailed species distributions, and distribution shifts between training and deployment data. To the best of our knowledge, the calibration of multi-label deep learning classifiers within the domain of bioacoustics has not yet been systematically assessed. We systematically benchmark the calibration of four state-of-the-art multi-label bird sound classifiers on the BirdSet benchmark, evaluating global, per-dataset, and per-class calibration using threshold-free calibration metrics (ECE, MCS) alongside discrimination metrics (cmAP). Model calibration varies significantly across datasets and classes. While Perch v2 and ConvNeXtBS show better global calibration, results vary between datasets. Both models indicate consistent underconfidence, while AudioProtoPNet and BirdMAE are mostly overconfident. Surprisingly, calibration seems to be better for less frequent classes. Using simple post hoc calibration methods we demonstrate a straightforward way to improve calibration. A small labelled calibration set is sufficient to significantly improve calibration with Platt scaling, while global calibration parameters suffer from dataset variability. Our findings highlight the importance of evaluating and improving uncertainty calibration in bioacoustic classifiers.
Figure 1: Calibration evaluation of AudioProtoPNet, BirdMAE, ConvNeXtBS and Perch v2 globally on all BirdSet test samples visualised in a reliability diagram. Mean predicted probabilities are shown on the x-axis, while empirical frequencies are shown on the y-axis. The miscalibration score (MCS) is outlined for each model.
@@ 1 INTRODUCTION
Human pressures are leading to declines in biodiversity, necessitating the development of measurement systems and effective analysis tools to monitor these trends (Schmeller et al.,, 2017; Gosselin and Callois,, 2018). Bioacoustics describes the study of sounds produced by organisms, offering a non-invasive measurement method of ecological habitats and their species (Laiolo,, 2010). Passive acoustic monitoring (PAM)has emerged as a powerful technique for continuous and non-invasive biodiversity assessments of animal populations in remote areas through low-cost recorders at scale (Sugai et al.,, 2019). Recent advances in deep learning (DL)have made it feasible to process large collections of PAM recordings and automatically detect species vocalisations, enabling applications such as population trend estimation (McGinn et al.,, 2023), impact assessments (Kershenbaum et al.,, 2025), and rapid alerts for protected or invasive species (Baumgartner et al.,, 2019). While the performance of DLmodels in bioacoustics has shown steady increases in previous years, generalisation to soundscape recordings with overlapping vocalisations, as collected in PAMscenarios, is still challenging (Stowell,, 2022; Rauch et al., 2025b, ; Schwinger et al.,, 2025). In practice, decisions based on classifier outputs rarely stand alone: ecologists, regulators, and land managers need to understand when to rely on model outputs, and when to defer to human review or collect more data. Variability in model calibration across domains calls into question the validity of fixed thresholds when applied across different deployment settings (Navine et al.,, 2024). Reliable decision-making therefore requires not only high discrimination (e.g., AUROC, cmAP scores) but also well-calibrated predictive probabilities that honestly reflect uncertainty (Guo et al.,, 2017).
@@
Calibration is especially challenging in bird sound classification as the task is inherently multi-label due to overlapping vocalisations. In addition, soundscapes vary widely across habitats and devices, and species occurrences follow a long tail, with rare species being of greatest conservation concern (Stowell,, 2022). Distribution shift between training data (often directional focal recordings) and deployment data (omnidirectionally recorded soundscapes) can further degrade reliability (Rauch et al., 2025b, ). In these settings, overconfident models can bias occupancy predictions, inflate apparent detection rates, and mislead thresholding and decision making, leading to samples not flagged for human evaluation. Conversely, underconfidence reduces the utility of scores for prioritising expert review. Reliable uncertainty estimates are therefore critical to support transparent and accountable use of PAM in environmental decision-making.
@@
Despite growing interest, the calibration behaviour of multi-label audio classifiers remains underexplored. Many studies prioritise accuracy or ranking metrics, while probability quality is reported sparsely or via single-number summaries that may obscure class-specific effects (van Merriënboer et al.,, 2025; Kahl et al.,, 2021). Moreover, commonly used post hoc calibration methods, which improve model calibration without retraining, are not uniformly evaluated in this domain. Based on this motivation, our contributions can be summarised as follows:
@@
Contributions C1. Benchmark calibration of four state-of-the-art multi-label bird sound classifiers on BirdSet, reporting global, per-dataset and per-class calibration with threshold-free metrics, revealing that per-domain (e.g., dataset, class, etc.) calibration is highly variable. C2. Investigate post hoc strategies (temperature and Platt scaling) showing that Platt scaling with 10 minutes of test data can significantly improve calibration. C3. Summarise implications to give practitioners concrete suggestions for deployment-specific calibration evaluation and improvement. Code is available online [^1] https://github.com/DBD-research-group/UncertainBird.
@@ 2.1 Uncertainty and Calibration
We follow the common decomposition into (i) aleatoric uncertainty inherent in the data (e.g., noise, wind, overlapping vocalisations, distant calls) and (ii) epistemic uncertainty due to a lack of model knowledge (e.g., rare species, novel soundscapes, or data-scarce regions) (Rewicki and Gawlikowski,, 2022). While we are primarily motivated by epistemic uncertainty, the calibration methods we employ are general-purpose post hoc techniques; we assume that acquiring more data is not possible.
@@
Calibration methods can be broadly divided into post hoc and ad hoc approaches. Post hoc methods operate on a trained model’s outputs, learning a mapping from uncalibrated logits to calibrated probabilities without modifying the model weights. Temperature scaling (TS) (Guo et al.,, 2017), Platt scaling (PS) (Platt,, 1999), isotonic regression (Niculescu-Mizil and Caruana,, 2005), and vector scaling (Kull et al.,, 2019) are widely used post hoc techniques. These methods are attractive due to their simplicity and low computational overhead. Bayesian deep learning methods, such as Monte Carlo dropout (Gal and Ghahramani,, 2016), deep ensembles (Lakshminarayanan et al.,, 2017), and variational inference (Blundell et al.,, 2015; Rizos et al.,, 2024), provide principled uncertainty estimates but add computational overhead.
Ad hoc calibration, in contrast, integrates calibration objectives directly into model training. This includes approaches such as label smoothing (Müller et al.,, 2019), focal loss (Lin et al.,, 2017), and mixup (Zhang et al.,, 2018), which implicitly encourage calibrated predictions by modifying the loss. More recently, spectral-normalised neural Gaussian processes (Liu et al.,, 2020) have been proposed to enhance distance awareness in DL, potentially improving calibration. Ad hoc methods can yield better calibration but often require careful tuning and may increase training complexity.
@@ 2.2 Platt and Temperature Scaling
Modern bioacoustic models typically formulate animal sound detection as a multi-label classification problem (Rauch et al., 2025b, ; van Merriënboer et al.,, 2025; Heinrich et al.,, 2025). For each class $c\in\{1,\dots,C\}$, the model produces a logit $z_{c}\in\mathbb{R}$, which is mapped to a confidence score $\hat{p}_{c}\in[0,1]$ via the sigmoid function, i.e., $\hat{p}_{c}=\sigma(z_{c})$. The resulting confidence scores can be calibrated post hoc without retraining the base model (Platt,, 1999; Niculescu-Mizil and Caruana,, 2005; Guo et al.,, 2017). A widely used method is TS (Guo et al.,, 2017), which rescales the logits by a positive temperature parameter $T\in\mathbb{R}_{>0}$ and then re-applies the sigmoid:
@@
The parameter $T$ controls the sharpness of the resulting distribution of calibrated confidence scores, see Figure 2(a). Setting $T=1$ recovers the uncalibrated model. For $T>1$, confidences are smoothed. Values below $0.5$ move up toward $0.5$, and values above $0.5$ move down toward $0.5$. For $T<1$, the distribution is sharpened, pushing values below $0.5$ toward 0 and values above $0.5$ toward $1$. If the same temperature parameter is used for all classes $c$, TSpreserves the monotonic ordering of the model’s predicted confidence scores (Guo et al.,, 2017; Tomani et al.,, 2022). Because $\sigma(0)=0.5$, confidence scores exactly at $0.5$ remain unchanged under TS. As a consequence, purely systematic offset errors (e.g., when the reliability curve (Murphy and Winkler,, 1977) lies consistently above or below the diagonal) cannot be corrected in this way (Alexandari et al.,, 2020). If offset errors need to be corrected, PS (Platt,, 1999) provides a useful generalisation of TS:
@@
In addition to the scaling parameter $T$, a bias $b\in\mathbb{R}$ is learned that shifts the entire sigmoid curve up or down, thereby changing the point at which a confidence score of $0.5$ is achieved (see Figures 2(b) and 2(c)). Positive biases ($b>0$) increase confidence scores, while negative biases ($b<0$) decrease them. Both TSand Platt scaling can also be parameterised on a per-class basis, using class-specific parameters $T_{c}$ and $b_{c}$ to compensate for class-dependent underconfidence and overconfidence (Alexandari et al.,, 2020; Frenkel and Goldberger,, 2021). In practice, the parameters are estimated by minimising the negative log-likelihood (NLL)of the true labels on a held-out validation set (Lin et al.,, 2007; Guo et al.,, 2017). The resulting optimisation problems are convex and involve only a few parameters to learn (Lin et al.,, 2007; Guo et al.,, 2017), making them well suited even when only limited data are available for calibration (Niculescu-Mizil and Caruana,, 2005). After fitting, the learned parameters are applied to all future predictions without modifying the base model.
@@ 2.3 Related work
2.3.1 Uncertainty calibration in multi-label classification
Multi-label classification presents unique calibration challenges due to label imbalance and the presence of co-occurring classes. Recent research has highlighted that standard deep neural networks, when trained for multi-label tasks, often produce poorly calibrated probability estimates, especially in the presence of severe class imbalance and large label spaces. Cheng and Vasconcelos, (2024) demonstrate that popular asymmetric loss functions, while effective for improving accuracy in unbalanced datasets, do not guarantee well-calibrated outputs. They attribute this to the lack of the strictly proper property in these losses and propose a new Strictly Proper Asymmetric loss, along with a label pair regulariser, to improve calibration without sacrificing accuracy. Their experiments show significant reductions in calibration error across diverse multi-label datasets.
In the context of extreme multi-label classification (XMLC), where the number of possible labels can reach into the millions, Ullah et al., (2024) systematically evaluate the calibration of a wide range of XMLC models. They find that naive application of standard metrics such as ECE can be misleading in long-tailed label distributions. To address this, they introduce calibration@k metrics (e.g., ECE@k), which focus on the top-k predicted labels, providing a more meaningful assessment of calibration in XMLC scenarios. Furthermore, they show that post hoc calibration using isotonic regression can substantially improve reliability without degrading predictive performance, making it a practical tool for practitioners. In bioacoustics, the use of pretrained models (e.g. Perch v2 (van Merriënboer et al.,, 2025), BirdNET (Kahl et al.,, 2021)) is increasingly common due to a lack of labelled in-domain data. At the same time, the number of classes for which these models compute predictions is increasing, with models such as Perch v2 having almost 15,000 classes available.
@@ 2.3.2 Uncertainty Calibration in Audio Classification
Uncertainty calibration for deep audio classifiers is an emerging research area, motivated by the need for reliable confidence estimates in practical applications such as multimedia retrieval, urban sound monitoring, and bioacoustic surveys. While deep neural networks have achieved state-of-the-art performance in audio classification tasks (Chen et al.,, 2023), model calibration is not well investigated yet.
Ye et al., (2022) provide one of the first systematic investigations of calibration methods for deep audio classifiers. They benchmark the calibration methods Monte Carlo Dropout, deep ensembles, focal loss, and spectral-normalised neural Gaussian processes (SNGP) on environmental sound and music genre classification datasets. Their results indicate that uncalibrated models are frequently overconfident, and that SNGP in particular achieves strong calibration and computational efficiency. Notably, while focal loss can improve calibration, it may degrade out-of-distribution detection performance, highlighting trade-offs between calibration and other uncertainty-related objectives.
In the context of bioacoustics, calibration is further complicated by the scarcity of strongly labelled data and the prevalence of rare species (Stowell,, 2022). Rizos et al., (2024) propose a Bayesian variational method which propagates uncertainty (prediction and variance) using a Bayesian neural network and is used for post hoc label smoothing.
To the best of our knowledge, our work is the first to benchmark calibration of state-of-the-art bioacoustic classifiers.
Table 1: Overview of the different datasets in BirdSet (Rauch et al., 2025b, ). J refers to the class imbalance index.
mainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmainmain Set ||Annotations|| ||Duration|| $\boldsymbol{J}$ $\boldsymbol{C}$ Train XCL 712,433  7,200h 0.85 9,734 Val POW 16,052 6.3h 0.66 48 Test PER 14,798 21h 0.78 132 NES 6,952 34h 0.76 89 UHH 59,583 50.9h 0.64 25 HSN 10,296 16.7h 0.54 21 NBP 5,493 0.8h 0.92 51 SSW 50,760 285h 0.77 81 SNE 20,147 33h 0.70 56